Write and solve an equation to determine the value of a $8000 investment if it is invested for 3 years at 3.5% compunded annually.
Answer by nyc_function(2741)

We use A = P times (1 + r)^(number of years)
A = P(1 + r)^n
A = 8000(1 + 0.035)^3
A = 8000(1.035)^3
A = 8000(1.108717875)
A = $8869.743
You can round answer off to the nearest hundreths. In that case, A = $8869.74